As a commodity, cocoa is traded in dollars per metric ton (tonne). On one day of trading cocoa futures saw normally distributed trading prices having a mean of $2,562/ tonne with a standard deviation of $12/ tonne. Determine the trading prices above which only 16% of the trading prices would lie.
Answer #1

solution:- Given that, ll=2562 6 -12 using standard normal table PC 2 > 2) = 161. 1-PC2L2) = 0.16 eczLz) = 1-0.16 =0.84 PCzlo